OpenStudy (anonymous):

what is cos 45 besides square root 2/2

jimthompson5910 (jim_thompson5910):

Sometimes you may see it as \[\Large \cos(45^{\circ}) = \frac{1}{\sqrt{2}}\] but many books rationalize the denominator

OpenStudy (anonymous):

or even the square root of one half same thing

OpenStudy (anonymous):

Thanks, both of you!